Good Morning,
PLEASE Help learn how to solve the below methods.
1.Solve using the elimination method.
5r - 6s = -25
6r - 5s = 31
2. Solve the system of equation using the substitution methods.
2x + 3y = -3
2x -y = 6
Thanking you in advance Betty

Answer by psbhowmick(878)

5r - 6s = -25 ------ (1)
6r - 5s = 31 ------ (2)
Multiply (2) with 5 and (1) with 6 and then subtracting (2) from (1)
6(5r - 6s) - 5(6r - 5s) = 6(-25) - 5(31)
Can take this forward from here?